OpenStudy (milo123):

If y = 5x − 4, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(0, −4), (2, 6), (4, 20)} {(0, −4), (2, 6), (4, 16)} {(0, 4), (2, 6), (4, 16)} {(0, 4), (2, −6), (4, 20)}

OpenStudy (gunsodo):

Let x = 0, 2, and 4 And then what choice has the correct y-values, that's correct one

OpenStudy (caozeyuan):

for example, if x is 0, the f=5*0-4=-4

OpenStudy (eliesaab):

If x=0, then y=-4, so c and d are not good answers

OpenStudy (eliesaab):

x=2, then y[?

OpenStudy (eliesaab):

y?

OpenStudy (eliesaab):

2(5)-4=6

OpenStudy (eliesaab):

if x=4, y=5(4)-4=16

OpenStudy (eliesaab):

So a is out. What is left?

OpenStudy (milo123):

B?

OpenStudy (eliesaab):

Yes